# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ B.2 Proofs of Section 4
$g(U+V)-g(U)=\mathcal{C}_{\textup{pos}}^{1/2}((U+V)(U+V)^{*}-UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(UU^{*}+VV^{*}+UV^{*}+VU^{*}).$
@@
$\lVert g(U+V)-g(U)-\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert\leq\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert VV^{*}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert V\rVert^{2}$
Dividing by $\lVert V\rVert$ and letting $\lVert V\rVert\rightarrow 0$, this shows that $g$ is differentiable with the correct derivative.
To show differentiability of $F_{f}$, let $x=(x_{i})_{i}\in\ell^{2}((-1,\infty))$, $y=(y_{i})_{i}\in\ell^{2}(\mathbb{R})$ and define $c_{x}\in\ell^{2}(\mathbb{R})$ by $(c_{x})_{i}=f^{\prime}(x_{i})$. By the assumption $f\in\mathscr{F}$, $f^{\prime}$ is Lipschitz continuous in some neighbourhood $(-a,a)$ of 00, with $a>0$ and with Lipschitz constant $M_{0}$. Let us take $N_{x}$ so large that $\lvert x_{i}\rvert<a/2$ for $i>N_{x}$. Let $\varepsilon>0$ arbitrary. By differentiability of $f$, we can choose $\delta_{x,\varepsilon}>0$ such that $x_{i}+z\in(-1,\infty)$ and $\lvert(f(x_{i}+z)-f(x_{i})-f^{\prime}(x_{i})z)\rvert<\varepsilon\lvert z\rvert$ for $\lvert z\rvert<\delta_{x,\varepsilon}$ and $i=1,\ldots,N_{x}$. We then have for $\lVert y\rVert<\min(\delta_{x,\varepsilon},a/2,\varepsilon)$,
@@
See 4.12
Proof of Proposition 4.12.
Any statement regarding differentiability in this proof should be understood as Fréchet differentiability. Let us write Sym($m$) for the symmetric matrices on $\mathbb{R}^{m}$ endowed with the trace-inner product: $\langle A,B\rangle_{\text{Sym(}m\text{)}}=\operatorname{tr}\,\left({BA}\right)$ for symmetric matrices $A$ and $B$ on $\mathbb{R}^{m}$. The result of [19, Theorem 1.1] is a statment on Sym($m$), and we therefore identify Sym($m$) with $L_{2}(\mathcal{Z})_{\mathbb{R}}$.
Let $(\varphi_{i})_{i}$ the standard basis of $\mathbb{R}^{m}$. Define $\Phi:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\text{Sym(}m\text{)}$ by $\Phi(X)=\sum_{i,j=1}^{m}\langle Xe_{j},e_{i}\rangle\varphi_{i}\otimes\varphi_{j}$ and let $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. Then $\Phi$ is an isomorphism of Hilbert spaces and by linearity of $\Phi$ we have $\Phi^{\prime}(X)(X_{2})=\sum_{i,j=1}^{m}\langle X_{2}e_{j},e_{i}\rangle\varphi_{i}\otimes\varphi_{j}$ for all $X_{2}\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. Furthermore, $\Lambda^{m}\circ\Phi^{-1}$ is the eigenvalue map on Sym($m$), where the eigenvalues are ordered in a nonincreasing way. We note that $\Phi(X)=\sum_{i=1}^{m}(\Lambda^{m}\circ\Phi^{-1})_{i}(\Phi(X))\varphi_{i}\otimes\varphi_{i}$. By the chain rule, $G\circ\Lambda^{m}$ is differentiable in $X$ if and only $G\circ\Lambda^{m}\circ\Phi^{-1}$ is differentiable in $\Phi(X)$. By [19, Theorem 1.1], $G\circ\Lambda^{m}\circ\Phi^{-1}$ is differentiable in $\Phi(X)$ if and only if $G$ is differentiable in $\Lambda^{m}\circ\Phi(\Phi(X))=\Lambda^{m}(X)$, and the derivative is given by
$(G\circ\Lambda^{m}\circ\Phi^{-1})^{\prime}(\Phi(X))=\sum_{i=1}^{m}G^{\prime}\left(\Lambda^{m}\circ\Phi^{-1}(\Phi(X))\right)_{i}\varphi_{i}\otimes\varphi_{i}=\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))_{i}\varphi_{i}\otimes\varphi_{i}.$
Thus, $G\circ\Lambda^{m}$ is differentiable in $X$ if and only if $G$ is differentiable in $\Lambda^{m}(X)$. Another application of the chain rule, the expression for $\Phi^{\prime}$ and the previous equation then finish the proof by showing that
$\langle(G\circ\Lambda^{m})^{\prime}(X),X_{2}\rangle_{L_{2}(\mathcal{X})}$ $=\langle(G\circ\Lambda^{m}\circ\Phi^{-1}\circ\Phi)^{\prime}(X),X_{2}\rangle_{L_{2}(\mathcal{Z})}$
$=\Big{\langle}(G\circ\Lambda^{m}\circ\Phi^{-1})^{\prime}(\Phi(X)),\Phi^{\prime}(X)(X_{2})\Big{\rangle}_{\text{Sym(}m\text{)}}$
$=\Big{\langle}\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))_{i}\varphi_{i}\otimes\varphi_{i},\sum_{k,j=1}^{m}\langle X_{2}e_{j},e_{k}\rangle_{L_{2}(\mathcal{H})}\varphi_{k}\otimes\varphi_{j}\Big{\rangle}_{\text{Sym(}m\text{)}}$
$=\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))_{i}\langle X_{2}e_{i},e_{i}\rangle_{L_{2}(\mathcal{H})}$
$=\langle\sum_{i=1}^{m}G^{\prime}(\Lambda^{m}(X))e_{i}\otimes e_{i},X_{2}\rangle_{L_{2}(\mathcal{Z})}.$
@@
Any statement regarding differentiability in this proof should be understood as Fréchet differentiability. For $X\in\mathcal{W}$, we have $\ker{X}^{\perp}=\ker{X^{*}}^{\perp}=\overline{\operatorname{ran}{X}}=\operatorname{ran}{X}$ for $X\in\mathcal{W}$ by Lemma A.7. Let $m\coloneqq\dim{\mathcal{Z}}$ and extend $(e_{i})_{i}$ to an ONB of $\mathcal{Z}$. Note that the $m^{2}$ operators $e_{i}\otimes e_{j}$, $i,j\leq m$, span the space $\mathcal{W}$. Therefore, $\dim{\mathcal{W}}=m^{2}<\infty$. Because finite-dimensional spaces are closed, we can define $P_{\mathcal{Z}}:\mathcal{H}\rightarrow\mathcal{Z}$, the orthogonal projector onto $\mathcal{Z}$. Furthermore, we let $P_{m}:\ell^{2}(\mathbb{R})\rightarrow\mathbb{R}^{m}$ be the orthogonal projector onto the first $m$-coordinates of an $\ell^{2}$ sequence. Thus, $P_{m}^{*}$ is the natural embedding of $\mathbb{R}^{m}$ into $\ell^{2}(\mathbb{R})$. Since Proposition 4.12 is a statement on $L_{2}$ spaces of finite dimension, we identify $\mathcal{W}$ with $L_{2}(\mathcal{Z})_{\mathbb{R}}$ via $\Phi:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow L_{2}(\mathcal{Z})_{\mathbb{R}}$ by $\Phi(X)=P_{\mathcal{Z}}{\left.\kern-1.2ptX\vphantom{\big{|}}\right|_{Z}}$. Now, $\Phi$ is linear and surjective and $\lVert\Phi\rVert\leq 1$. For $X\in\mathcal{W}$, $\operatorname{ran}{X}\subset\mathcal{Z}$ and $\ker{X}^{\perp}=\operatorname{ran}{X}$, hence $\lVert\Phi(X)\rVert=\lVert{\left.\kern-1.2ptX\vphantom{\big{|}}\right|_{\mathcal{Z}}}\rVert=\lVert X\rVert$ and $\Phi$ is an isometry on $\mathcal{W}$.
We first prove the result for the case in which $\Lambda$ assigns a nonincreasing order to eigenvalues of nonnegative operators in $L_{2}(\mathcal{Z})_{\mathbb{R}}$. With $\Lambda^{m}$ defined before Proposition 4.12, we then have $\Lambda(X)_{i}=\Lambda^{m}(\Phi(X))_{i}$ for all $i\leq m$ and $X\in\mathcal{W}$. This implies that
$P_{m}\Lambda(X)$ $=\Lambda^{m}(\Phi(X)),\quad X\in\mathcal{W},$
$\Lambda(X)$ $=P_{m}^{*}\Lambda^{m}(\Phi(X)),\quad X\in\mathcal{W},$ (39)
since any $X\in\mathcal{W}$ has at most $m$ nonzero eigenvalues. With $F^{m}_{f}\coloneqq FP_{m}^{*}$, we also have
$(F\circ\Lambda)(X)=(F^{m}_{f}\circ\Lambda^{m})(\Phi(X)),\quad X\in\mathcal{W}.$ (40)
By the hypothesis on $F$, $F$ is differentiable in $\Lambda(X)$ for $X\in\mathcal{W}$. Furthermore, $P_{m}^{*}$ is linear, hence differentiable, and $(P_{m}^{*})^{\prime}(x)(y)=P_{m}^{*}y$ for $x,y\in\mathbb{R}^{m}$. Then, by (39) and the chain rule, the composition $F^{m}_{f}=F\circ P_{m}^{*}$ is differentiable in $\Lambda^{m}(\Phi(X))$, and by (39) it holds for any $y\in\mathbb{R}^{m}$,
$\langle(F^{m}_{f})^{\prime}(\Lambda^{m}(\Phi(X))),y\rangle=\langle F^{\prime}(P_{m}^{*}\Lambda^{m}(\Phi(X))),P_{m}^{*}y\rangle=\langle P_{m}F^{\prime}(\Lambda(X)),y\rangle,$
i.e.
$(F^{m})^{\prime}(\Lambda^{m}(\Phi(X)))=P_{m}F^{\prime}(\Lambda(X))\in\mathbb{R}^{m}.$ (41)
We may now apply Proposition 4.12 to conclude that $F^{m}\circ\Lambda^{m}$ is differentiable in $\Phi(X)$. By the fact that $e_{i}\in\operatorname{ran}{Z}$ and by the stated diagonalisation of $X$, we have $\Phi(X)=\sum_{i=1}^{m}\Lambda^{m}(\Phi(X))_{i}e_{i}\otimes e_{i}$, where the rank-1 operators $e_{i}\otimes e_{i}$ are now understood to act only on $\mathcal{Z}$. With Proposition 4.12 we thus also obtain the expression for the derivative
@@
$(F^{m}\circ\Lambda^{m})^{\prime}(\Phi(X))=\sum_{i=1}^{m}(F^{m})^{\prime}(\Lambda^{m}(\Phi(X)))_{i}e_{i}\otimes e_{i}=\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{Z})_{\mathbb{R}},$ (42)
@@
where for second equation we use (41), the definition of $P_{m}$ and the fact that the sums extend only upto $m$. Another application of the chain rule to (40) now shows that $F\circ\Lambda$ is differentiable in $X$ and, because $\Phi^{\prime}(X)(Y)=\Phi(Y)$ for $Y\in\mathcal{W}$, that
@@
$\langle(F\circ\Lambda)^{\prime}(X),Y\rangle_{L_{2}(\mathcal{H})}$ $=\langle(F^{m}\circ\Lambda^{m})^{\prime}(\Phi(X)),\Phi(Y)\rangle_{L_{2}(\mathcal{Z})}=\Big{\langle}\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},\Phi(Y)\Big{\rangle}_{L_{2}(\mathcal{Z})}$
$=\Big{\langle}\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},Y\Big{\rangle}_{L_{2}(\mathcal{\mathcal{H}})}=\Big{\langle}\sum_{i=1}^{\infty}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},Y\Big{\rangle}_{L_{2}(\mathcal{\mathcal{H}})},$
where in the third step we use $Y\in\mathcal{W}$ and again interpret $e_{i}\otimes e_{i}$ as acting on all of $\mathcal{H}$, and in the last step we use that the linear span of $(e_{i})_{i=1}^{m}$ contains $\operatorname{ran}{Y}\subset\mathcal{Z}$. Since $Y\in\mathcal{W}$ is arbitrary, this shows that ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is differentiable in $X$.
Finally, let us denote by $\tilde{\Lambda}$ an eigenvalue map on $L_{2}(\mathcal{H})_{\mathbb{R}}$ which can assign any fixed but arbitrary ordering on the eigenvalues. Let $X\in\mathcal{W}$, $X=\sum_{i=1}\tilde{\Lambda}(X)_{i}e_{i}\otimes e_{i}$ be given and assume that $F$ is differentiable in $\tilde{\Lambda}(X)$. Given $X$, there exists a permutation $\alpha:\mathbb{N}\rightarrow\mathbb{N}$ such that $\tilde{\Lambda}(X)_{i}=\Lambda(X)_{\alpha(i)}$ for all $i\in\mathbb{N}$. Let $P_{\alpha}:\ell^{2}(\mathbb{R})\rightarrow\ell^{2}(\mathbb{R})$ denote the permutation operator $(P_{\alpha}x)_{i}=(Px)_{\alpha(i)}$, $i\in\mathbb{N}$, so $P_{\alpha}^{*}=P_{\alpha}^{-1}$. Then $\tilde{\Lambda}(X)=(P_{\alpha}\Lambda)(X)$ and $X=\sum_{i}\Lambda_{\alpha(i)}(X)e_{i}\otimes e_{i}=\sum_{i}\Lambda(X)_{i}e_{\alpha^{-1}(i)}\otimes e_{\alpha^{-1}(i)}.$ By the previous part of the proof, ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is differentiable in $X$. Because $F\circ\tilde{\Lambda}=F\circ\Lambda$ by symmetry of $F$, differentiability of ${\left.\kern-1.2pt(F\circ\tilde{\Lambda})\vphantom{\big{|}}\right|_{\mathcal{W}}}$ in $X$ follows, and
$(F\circ\tilde{\Lambda})^{\prime}(X)=(F\circ\Lambda)^{\prime}(X)=\sum_{i}F_{i}^{\prime}(\Lambda(X))e_{\alpha^{-1}(i)}\otimes e_{\alpha^{-1}(i)}.$ (43)
Since $F$ is symmetric, $F\circ P_{\alpha}=F$. Hence, for $x\in\ell^{2}(\mathbb{R})$,
$F^{\prime}(x)=(F\circ P_{\alpha})^{\prime}(x)=P_{\alpha}^{*}F^{\prime}(P_{\alpha}x)=P_{\alpha}^{-1}F^{\prime}(P_{\alpha}x).$
Thus, $F_{i}^{\prime}(\Lambda(X))=F_{\alpha^{-1}(i)}^{\prime}(P_{\alpha}\Lambda(X))=F_{\alpha^{-1}(i)}^{\prime}(\tilde{\Lambda}(X))$. From (43) it now follows that
$(F\circ\tilde{\Lambda})^{\prime}(X)=\sum_{i}F_{\alpha^{-1}(i)}^{\prime}(\tilde{\Lambda}(X))e_{\alpha^{-1}(i)}\otimes e_{\alpha^{-1}(i)}=\sum_{i}F_{i}^{\prime}(\tilde{\Lambda}(X))e_{i}\otimes e_{i}.$
See 4.15

### v3

@@ B.2 Proofs of Section 4
$g(U+V)-g(U)=\mathcal{C}_{\textup{pos}}^{1/2}((U+V)(U+V)^{*}-UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*}+UV^{*}+VU^{*}).$
@@
$\lVert g(U+V)-g(U)-\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert\leq\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert VV^{*}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert V\rVert^{2}.$
Dividing by $\lVert V\rVert$ and letting $\lVert V\rVert\rightarrow 0$, this shows that $g$ is differentiable and has the stated derivative.
To show differentiability of $F_{f}$, let $x=(x_{i})_{i}\in\ell^{2}((-1,\infty))$, $y=(y_{i})_{i}\in\ell^{2}(\mathbb{R})$ and define $c_{x}\in\ell^{2}(\mathbb{R})$ by $(c_{x})_{i}=f^{\prime}(x_{i})$. By the assumption $f\in\mathscr{F}$, $f^{\prime}$ is Lipschitz continuous in some neighbourhood $(-a,a)$ of 00, with $a>0$ and with Lipschitz constant $M_{0}$. Let us take $N_{x}$ so large that $\lvert x_{i}\rvert<a/2$ for $i>N_{x}$. Let $\varepsilon>0$ be arbitrary. By differentiability of $f$, we can choose $\delta_{x,\varepsilon}>0$ such that $x_{i}+z\in(-1,\infty)$ and $\lvert(f(x_{i}+z)-f(x_{i})-f^{\prime}(x_{i})z)\rvert<\varepsilon\lvert z\rvert$ for $\lvert z\rvert<\delta_{x,\varepsilon}$ and $i=1,\ldots,N_{x}$. We then have for $\lVert y\rVert<\min(\delta_{x,\varepsilon},a/2,\varepsilon)$,
@@
We need to relate the Fréchet differentiability of the composition $\mathcal{G}\circ\Lambda^{m}$, being the composition of the eigenvalue map $\Lambda^{m}$ on $L_{2}(\mathcal{Z})_{\mathbb{R}}$ as defined in Section 1.5 and a symmetric function $\mathcal{G}$, to Fréchet differentiability of $\mathcal{G}$ itself. To do so, we use [20, Theorem 1.1], which states this for the case that $\Lambda^{m}$ is defined on the space of symmetric matrices instead of $L_{2}(\mathcal{Z})$. Therefore, we identify this space of symmetric matrices with $L_{2}(\mathcal{Z})_{\mathbb{R}}$. The details of this identification are described in the following.
Any statement regarding differentiability in this proof should be understood as Fréchet differentiability. Let us write Sym($m$) for the symmetric matrices on $\mathbb{R}^{m}$ endowed with the trace-inner product: $\langle A,B\rangle_{\text{Sym(}m\text{)}}=\operatorname{tr}\,\left({BA}\right)$ for symmetric matrices $A,B\in\mathbb{R}^{m\times m}$. Let $(\varphi_{i})_{i}$ be the standard basis of $\mathbb{R}^{m}$. Define $\Phi:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\text{Sym(}m\text{)}$ by $\Phi(X)=\sum_{i,j=1}^{m}\langle Xe_{j},e_{i}\rangle\varphi_{i}\otimes\varphi_{j}$ and let $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. Then $\Phi$ is an isomorphism of Hilbert spaces and by linearity of $\Phi$ we have $\Phi^{\prime}(X)(X_{2})=\sum_{i,j=1}^{m}\langle X_{2}e_{j},e_{i}\rangle\varphi_{i}\otimes\varphi_{j}$ for all $X_{2}\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. Furthermore, $\Lambda^{m}\circ\Phi^{-1}$ is the eigenvalue map on Sym($m$), where the eigenvalues are ordered in a nonincreasing way. We note that $\Phi(X)=\sum_{i=1}^{m}(\Lambda^{m}\circ\Phi^{-1})_{i}(\Phi(X))\varphi_{i}\otimes\varphi_{i}$. Because $\Omega$ and $\mathcal{G}$ are symmetric by hypothesis, we may apply [20, Theorem 1.1], which states that $\mathcal{G}\circ(\Lambda^{m}\circ\Phi^{-1})$ is differentiable in $\Phi(X)$ if and only if $\mathcal{G}$ is differentiable in $\Lambda^{m}\circ\Phi^{-1}(\Phi(X))=\Lambda^{m}(X)$, in which case the derivative is given by
@@
$(\mathcal{G}\circ\Lambda^{m}\circ\Phi^{-1})^{\prime}(\Phi(X))=\sum_{i=1}^{m}\mathcal{G}^{\prime}\left(\Lambda^{m}\circ\Phi^{-1}(\Phi(X))\right)_{i}\varphi_{i}\otimes\varphi_{i}=\sum_{i=1}^{m}\mathcal{G}^{\prime}(\Lambda^{m}(X))_{i}\varphi_{i}\otimes\varphi_{i}.$
@@
By the chain rule, $\mathcal{G}\circ\Lambda^{m}$ is differentiable in $X$ if and only if $\mathcal{G}\circ\Lambda^{m}\circ\Phi^{-1}$ is differentiable in $\Phi(X)$. Thus, by the above display, $\mathcal{G}\circ\Lambda^{m}$ is differentiable in $X$ if and only if $\mathcal{G}$ is differentiable in $\Lambda^{m}(X)$. Another application of the chain rule, the expression for $\Phi^{\prime}$ and the previous equation then finish the proof by showing that
$\langle(\mathcal{G}\circ\Lambda^{m})^{\prime}(X),X_{2}\rangle_{L_{2}(\mathcal{Z})}$ $=\langle(\mathcal{G}\circ\Lambda^{m}\circ\Phi^{-1}\circ\Phi)^{\prime}(X),X_{2}\rangle_{L_{2}(\mathcal{Z})}$
$=\Big{\langle}(\mathcal{G}\circ\Lambda^{m}\circ\Phi^{-1})^{\prime}(\Phi(X)),\Phi^{\prime}(X)(X_{2})\Big{\rangle}_{\text{Sym(}m\text{)}}$
@@
$=\Big{\langle}\sum_{i=1}^{m}\mathcal{G}^{\prime}(\Lambda^{m}(X))_{i}\varphi_{i}\otimes\varphi_{i},\sum_{k,j=1}^{m}\langle X_{2}e_{j},e_{k}\rangle\varphi_{k}\otimes\varphi_{j}\Big{\rangle}_{\text{Sym(}m\text{)}}$
$=\sum_{i=1}^{m}\mathcal{G}^{\prime}(\Lambda^{m}(X))_{i}\langle X_{2}e_{i},e_{i}\rangle$
$=\Big{\langle}\sum_{i=1}^{m}\mathcal{G}^{\prime}(\Lambda^{m}(X))e_{i}\otimes e_{i},X_{2}\Big{\rangle}_{L_{2}(\mathcal{Z})},$
for any $X_{2}\in L_{2}(\mathcal{Z})_{\mathbb{R}}$. ∎
See 4.16